OpenStudy (anonymous):

Write an equation of the line whose slope is 3 and which passes through the point (1,6)

OpenStudy (***[isuru]***):

the general form of an equation is y = mx + b where m = slope b = intercept now since u know that the slope is 3 u can rewrite this as y = 3x + b now to find be u can use the coordinate (1 , 6) if u plug them in to the equation 6 = 3 + b b= 3 so the equation of the line is y = 3x + 3 got it :) ?

OpenStudy (anonymous):

yes, Thank you

OpenStudy (***[isuru]***):

u r welcome !